# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ 4.1. Framed construction of $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$
Moreover, by the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the comultiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, and the definition of $\varphi$, we have
@@
$\varphi\pi_{\xi\odot\lambda}(b)=b_{1}\otimes b_{2}\ \mathrm{mod}\,v^{-1}\mathcal{L}_{\xi,\lambda}.$
@@
In particular, we have $\varphi\pi_{\xi\odot\lambda}(b)\in\mathcal{L}_{\xi,\lambda}$. Since $b\in\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}(\xi\odot\lambda)$, we know that $\pi_{\xi\odot\lambda}(b)$ is bar-invariant with respect to the bar-involution of $\Lambda_{\xi,\lambda}$, see Theorem 2.8. By Lemma 3.9, we have
$\Psi\varphi\pi_{\xi\odot\lambda}(b)=\varphi(\overline{\pi_{\xi\odot\lambda}(b)})=\varphi\pi_{\xi\odot\lambda}(b).$
@@
By the uniqueness of $b_{1}\diamondsuit b_{2}$, we have $\varphi\pi_{\xi\odot\lambda}(b)=b_{1}\diamondsuit b_{2}$. Since $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is an isomorphism, we conclude that the image of the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ of $\Lambda_{\xi,\lambda}$ is the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$. ∎
@@
By Theorem 2.9, we have $b_{1}\diamondsuit b_{2}\in b_{1}\otimes b_{2}+\sum v^{-1}\mathbb{Z}[v^{-1}]b^{\prime}_{1}\otimes b^{\prime}_{2}$. Suppose that $b_{1}\diamondsuit b_{2}=\varphi\pi_{\xi\odot\lambda}(b)$, where $b\in\tilde{\mathbf{B}}(\xi,\lambda)$. By the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the comultiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, and the definition of $\varphi$, we have $\varphi\pi_{\xi\odot\lambda}(b)\in\sum\mathbb{N}[v,v^{-1}]b^{\prime\prime}_{1}\otimes b^{\prime\prime}_{2}$, where the sum is taken over $b^{\prime\prime}_{1}\otimes b^{\prime\prime}_{2}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$. Since $v^{-1}\mathbb{Z}[v^{-1}]\cap\mathbb{N}[v,v^{-1}]=v^{-1}\mathbb{N}[v^{-1}]$, we have $b_{1}\diamondsuit b_{2}\in b_{1}\otimes b_{2}+\sum v^{-1}\mathbb{N}[v^{-1}]b^{\prime}_{1}\otimes b^{\prime}_{2}$. ∎
@@ 4.2. Example
We present an example for Theorem 4.7. Let $(I,\cdot)$ be the symmetric Cartan datum of type $A_{1}$ and $(Y,X,\langle,\rangle,...)$ be the simply connected root datum of type $(I,\cdot)$, that is, $I=\{i\},i\cdot i=2$ and $Y=\mathbb{Z}[I],X=\mathrm{Hom}(Y,\mathbb{Z})$ with the natural pairing $\langle,\rangle:Y\times X\rightarrow\mathbb{Z}$ and natural embeddings $I\hookrightarrow Y,I\hookrightarrow X$. Then the canonical basis of $\mathbf{f}$ is
@@
with the identification $\alpha_{k,l}=\beta_{k,l}$ for $k-l=m-n$, see [12, Section 6].
The framed Cartan datum $(\tilde{I},\cdot)$ is of type $A_{2}$. The canonical basis of $\tilde{\mathbf{f}}$ is
@@
with the identification $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}=\theta_{i}^{(l)}\theta_{i^{\prime}}^{(n)}\theta_{i}^{(m-k)}$ for $k-l=m-n$. Moreover,
@@
Notice that $\theta_{i^{\prime}}^{(t-s)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(n-t+s)}\in\tilde{\mathbf{B}}$. Comparing the coefficient of $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}$, we obtain $t=m-k$ or $t=n-l\leqslant m-k$. Hence we have $t=m-k\leqslant m$, and so $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}\in\tilde{\mathbf{B}}(m,n)$. Moreover, we have
@@
For any $-m+k\leqslant s<0$, by similarly argument of $t_{i}^{\sigma}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)})=m-k$, we have $t_{i}^{\sigma}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(l-s)}\theta_{i^{\prime}}^{(l)})=-s>0$, and so $\pi_{0\odot n}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(l-s)}\theta_{i^{\prime}}^{(l)})=0\in\Lambda_{0,\lambda}$. For any $0\leqslant s\leqslant l$, by [13, Lemma 42.1.2(d)], we have
@@ 4.3. Positivity
${{}_{i}r}(1)=0,\ {{}_{i}r}(\theta_{j})=\delta_{i,j},\ {{}_{i}r}(xy)={{}_{i}r}(x)y+v^{|x|\cdot i}x{{}_{i}r}(y),$
@@
(11) $r(x)=$ $\theta_{i}\otimes{{}_{i}r}(x)+\textrm{other homogeneous terms}$
$=$ $r_{i}(x)\otimes\theta_{i}+\textrm{other homogeneous terms}.$
@@
The first formula is well-known, for example, see the proof of [13, Lemma 3.2.2]. The second formula follows from [13, Lemma 1.2.14] and $\langle i,\xi\odot\lambda\rangle=\langle i,\xi\rangle$. ∎
@@
For any $i\in I$ and $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that ${{}_{i}r}(b)=\sum_{b^{\prime}\in\tilde{\mathbf{B}}}\sum_{n\in\mathbb{Z}}d_{b^{\prime},n}^{b}v^{n}b^{\prime}$ for any $b^{\prime}\in\tilde{\mathbf{B}}$ and $n\in\mathbb{Z}$, we have $d_{b^{\prime},n}\in\mathbb{N}$. Moreover, for any $b^{\prime}\in\tilde{\mathbf{B}}(\xi,\lambda)$ and $n\in\mathbb{Z}$ such that $d_{b^{\prime},n}^{b}\not=0$, we have
@@
The first statement follows from the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the comultiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, and (11). The second statement follows from $\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}(\xi\odot\lambda)$ and [13, Corollary 22.1.5]. ∎
@@
For any $b\in\tilde{\mathbf{B}}(\xi,\lambda)$, by the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the multiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, we have $\theta_{i}b\in\sum_{b^{\prime}\in\tilde{\mathbf{B}}}\mathbb{N}[v,v^{-1}]b^{\prime}$. By Corollary 4.5, we have
@@
The other statements follow from Theorem 3.8 and 4.7. ∎
@@ 4.4. Generalized case
Our method can be generalized to the case for the tensor product of several integrable highest weight modules by further enlarging the Cartan datum. For example, we sketch the construction for the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}\otimes\Lambda_{\zeta}$, where $\zeta$ is another dominant weight with respect to the root datum $(Y,X,\langle,\rangle,...)$. For the Cartan datum $(I,\cdot)$, we enlarge its framed Cartan datum $(\tilde{I},\cdot)$ to be the $2$-framed Cartan datum $(\tilde{I}^{2},\cdot)$, where $\tilde{I}^{2}=\tilde{I}\sqcup I^{\prime\prime}=\tilde{I}\sqcup\{i^{\prime\prime}\mid i\in I\}$ and
@@
Similarly, for the root datum $(Y,X,\langle,\rangle,...)$, we enlarge its framed root datum $(\tilde{Y},\tilde{X},\langle,\rangle,...)$ to be the $2$-framed root datum $(\tilde{Y}^{2},\tilde{X}^{2},\langle,\rangle,...)$. Then we have the corresponding algebras $\tilde{\mathbf{f}}^{2}$ with the canonical basis $\tilde{\mathbf{B}}^{2}$, and the quantized enveloping algebra $\tilde{\mathbf{U}}^{2}$. We define
@@
Then we have the corresponding Verma module $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$ and the simple highest weight module $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$ of $\tilde{\mathbf{U}}^{2}$ with the canonical basis $\tilde{\mathbf{B}}^{2}(\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2})$. We restrict them to be $\mathbf{U}$-modules, and can prove that the subspace $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{\mathbf{f}}^{2}$ spanned by $\{x\theta_{\zeta}y\theta_{\lambda}z\mid x,y,z\in\mathbf{f}\}$, where $\theta_{\zeta}=\prod_{i\in I}\theta_{i^{\prime\prime}}^{(\langle i,\zeta\rangle)}\in\tilde{\mathbf{f}}^{2}$, is a $\mathbf{U}$-submodule $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$. This is analogue to the submodule $\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{M}_{\xi\odot\lambda}$. We define the $\mathbf{U}$-module $\Lambda_{\xi,\lambda,\zeta}$ to be the image of $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ under the natural projection $\pi_{\xi\odot\lambda\odot\zeta}:\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}\rightarrow\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2}$. This is analogue to the module $\Lambda_{\xi,\lambda}$. Moreover, we can construct a subset $\tilde{\mathbf{B}}^{2}(\Lambda_{\xi,\lambda,\zeta})\subset\tilde{\mathbf{B}}^{2}(\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2})$, and prove that it is a basis of $\Lambda_{\xi,\lambda,\zeta}$. This is analogue to the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ of $\Lambda_{\xi,\lambda}$. We can construct a $\mathbf{U}$-module isomorphism $\varphi^{2}:\Lambda_{\xi,\lambda,\zeta}\rightarrow\Lambda_{\xi,\lambda}\otimes\Lambda_{\zeta}$ by using of the comultiplication of $\tilde{\mathbf{f}}^{2}$, and prove that the image of $\tilde{\mathbf{B}}^{2}(\Lambda_{\xi,\lambda,\zeta})$ under $\varphi^{2}$ coincides with the canonical basis of the tensor product $\Lambda_{\xi,\lambda}\otimes\Lambda_{\zeta}$. This is analogue to Theorem 3.8 and 4.7.

### v3

@@ 4.1. Framed construction of $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$
For any $b\in\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}$, suppose that $r(b)=\sum_{b^{\prime}_{1},b^{\prime}_{2}\in\tilde{\mathbf{B}}}c_{b^{\prime}_{1},b^{\prime}_{2}}b^{\prime}_{1}\otimes b^{\prime}_{2}$, where $c_{b^{\prime}_{1},b^{\prime}_{2}}\in\mathbb{N}[v,v^{-1}]$ by the positivity property of $\tilde{\mathbf{B}}$ with respect to the comultiplication. By the definition of $\varphi$, we have
$\varphi\pi_{\xi\odot\lambda}(b)=\sum c_{b^{\prime}_{1},b^{\prime}_{2}}{b^{\prime}_{2}}^{-}\eta_{\xi}\otimes\phi_{\lambda}^{-1}\pi_{0\odot\lambda}(b^{\prime}_{1}),$
where the summation is taken over $b^{\prime}_{1},b^{\prime}_{2}\in\tilde{\mathbf{B}}$ such that $|b^{\prime}_{1}|-|\theta_{\lambda}|\in\mathbb{N}[I]$. By the proof of Lemma 3.5 and Corollary 4.5, we have $\sum c_{b^{\prime}_{1},b^{\prime}_{2}}{b^{\prime}_{2}}^{-}\eta_{\xi}\otimes\pi_{0\odot\lambda}(b^{\prime}_{1})\in\Lambda_{\xi}\otimes\Lambda_{0,\lambda}$, where $\Lambda_{0,\lambda}=\pi_{0\odot\lambda}((\mathbf{f}\theta_{\lambda}\mathbf{f})_{0\odot\lambda})\subset\Lambda_{0\odot\lambda}$, and then ${b^{\prime}_{2}}^{-}\eta_{\xi}\in\mathbf{B}(\Lambda_{\xi}),b^{\prime}_{1}\in\mathbf{B}((\mathbf{f}\theta_{\lambda}\mathbf{f})_{0\odot\lambda})$. If $\pi_{0\odot\lambda}(b^{\prime}_{1})\not=0$, by definitions and (14), we have $b^{\prime}_{1}\in\tilde{\mathbf{B}}(0\odot\lambda)\subset\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}^{\sigma}$, and there exists a unique $b^{\prime\prime}_{1}\in\mathbf{B}(\lambda)\subset\mathbf{B}$ such that $b^{\prime}_{1}=b^{\prime\prime}_{1}\theta_{\lambda}$. Hence
@@
(7) $\varphi\pi_{\xi\odot\lambda}(b)=\sum c_{b^{\prime}_{1},b^{\prime}_{2}}{b^{\prime}_{2}}^{-}\eta_{\xi}\otimes{b^{\prime\prime}_{1}}^{-}\eta_{\lambda}\in\mathbb{N}[v,v^{-1}][\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})]\subset\mathcal{A}\otimes_{\mathbb{Z}[v^{-1}]}\mathcal{L}_{\xi,\lambda}.$
By Lemma 4.6, $(\varphi\pi_{\xi\odot\lambda}(b),\varphi\pi_{\xi\odot\lambda}(b))_{\xi,\lambda}\in 1+v^{-1}\mathbf{A}$. By Lemma 2.4, there exists $b_{1}^{-}\eta_{\xi}\otimes b_{2}^{-}\eta_{\lambda}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$ such that
$\varphi\pi_{\xi\odot\lambda}(b)=\pm b_{1}^{-}\eta_{\xi}\otimes b_{2}^{-}\eta_{\lambda}\ \mathrm{mod}\,v^{-1}L(\Lambda_{\xi}\otimes\Lambda_{\lambda}),$
where $L(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ is the $\mathbf{A}$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$. By (7) and $\mathcal{A}\cap\mathbf{A}=\mathbb{Z}[v^{-1}]$,
$\varphi\pi_{\xi\odot\lambda}(b)=\pm b_{1}^{-}\eta_{\xi}\otimes b_{2}^{-}\eta_{\lambda}\ \mathrm{mod}\,v^{-1}\mathcal{L}_{\xi,\lambda}=b_{1}^{-}\eta_{\xi}\otimes b_{2}^{-}\eta_{\lambda}\ \mathrm{mod}\,v^{-1}\mathcal{L}_{\xi,\lambda},$
@@
that is, $\varphi\pi_{\xi\odot\lambda}(b)-b_{1}^{-}\eta_{\xi}\otimes b_{2}^{-}\eta_{\lambda}\in v^{-1}\mathcal{L}_{\xi,\lambda}$. Since $b\in\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}(\xi\odot\lambda)$, we know that $\pi_{\xi\odot\lambda}(b)$ is bar-invariant with respect to the bar-involution of $\Lambda_{\xi,\lambda}$, see Theorem 2.8. By Lemma 3.9, we have
$\Psi\varphi\pi_{\xi\odot\lambda}(b)=\varphi\pi_{\xi\odot\lambda}(b).$
@@
By the uniqueness of $b_{1}^{-}\eta_{\xi}\diamondsuit b_{2}^{-}\eta_{\lambda}$, we have $\varphi\pi_{\xi\odot\lambda}(b)=b_{1}^{-}\eta_{\xi}\diamondsuit b_{2}^{-}\eta_{\lambda}$. Since $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is an isomorphism, we conclude that the image of the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ of $\Lambda_{\xi,\lambda}$ is the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$. ∎
@@
By Theorem 2.9, we have $b_{1}\diamondsuit b_{2}\in b_{1}\otimes b_{2}+\sum v^{-1}\mathbb{Z}[v^{-1}]b^{\prime}_{1}\otimes b^{\prime}_{2}$. Suppose that $b_{1}\diamondsuit b_{2}=\varphi\pi_{\xi\odot\lambda}(b)$, where $b\in\tilde{\mathbf{B}}(\xi,\lambda)$. By (7), we have $\varphi\pi_{\xi\odot\lambda}(b)\in\sum\mathbb{N}[v,v^{-1}]b^{\prime\prime}_{1}\otimes b^{\prime\prime}_{2}$, where the sum is taken over $b^{\prime\prime}_{1}\otimes b^{\prime\prime}_{2}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$. Since $v^{-1}\mathbb{Z}[v^{-1}]\cap\mathbb{N}[v,v^{-1}]=v^{-1}\mathbb{N}[v^{-1}]$, we have $b_{1}\diamondsuit b_{2}\in b_{1}\otimes b_{2}+\sum v^{-1}\mathbb{N}[v^{-1}]b^{\prime}_{1}\otimes b^{\prime}_{2}$. ∎
@@ 4.2. Example
[…]
@@
with the identification $\alpha_{k,l}=\beta_{k,l}$ for kl=mnk-l=m-n, see [12, Section 6].
[…]
@@
with the identification $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}=\theta_{i}^{(l)}\theta_{i^{\prime}}^{(n)}\theta_{i}^{(m-k)}$ for kl=mnk-l=m-n. Moreover,
@@
Notice that $\theta_{i^{\prime}}^{(t-s)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(n-t+s)}\in\tilde{\mathbf{B}}$. Comparing the coefficient of $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}$, we obtain t=mkt=m-k or $t=n-l\leqslant m-k$. Hence we have $t=m-k\leqslant m$, and so $\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)}\in\tilde{\mathbf{B}}(m,n)$. Moreover, we have
@@
For any $m+k\leqslant s<0$, by similarly argument of $t_{i}^{\sigma}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(m-k+l)}\theta_{i^{\prime}}^{(l)})=m-k$, we have $t_{i}^{\sigma}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(l-s)}\theta_{i^{\prime}}^{(l)})=-s>0$, and so $\pi_{0\odot n}(\theta_{i^{\prime}}^{(n-l)}\theta_{i}^{(l-s)}\theta_{i^{\prime}}^{(l)})=0\in\Lambda_{0,\lambda}$. For any $0\leqslant s\leqslant l$, by [13, Lemma 42.1.2(d)], we have
@@ 4.3. Positivity
(12) ${{}_{i}r}(1)=0,\ {{}_{i}r}(\theta_{j})=\delta_{i,j},\ {{}_{i}r}(xy)={{}_{i}r}(x)y+v^{|x|\cdot i}x{{}_{i}r}(y),$
@@
(13) $r(x)=\theta_{i}\otimes{{}_{i}r}(x)+\sum x_{1}\otimes x_{2}=r_{i}(x)\otimes\theta_{i}+\sum x^{\prime}_{1}\otimes x^{\prime}_{2},$
where the summations are finite, and $x_{1},x_{2},x^{\prime}_{1},x^{\prime}_{2}\in\mathbf{f}$ are homogeneous such that $|x_{1}|\not=i$ and $|x^{\prime}_{2}|\not=i$.
@@
Since $E_{i}x=E_{i}x^{-}1$ and $E_{i}1=0$ in $\tilde{M}_{\xi\odot\lambda}$, by [13, Proposition 3.1.6], we have
$E_{i}x=\frac{K_{i}({{}_{i}r(x)})^{-}1-(r_{i}(x))^{-}K_{-i}1}{v-v^{-1}}=\frac{v^{\langle i,\xi\odot\lambda-|x|\rangle+2}{{}_{i}r}(x)-v^{-\langle i,\xi\odot\lambda\rangle}r_{i}(x)}{v-v^{-1}}.$
The second formula follows from [13, Lemma 1.2.14] and $\langle i,\xi\odot\lambda\rangle=\langle i,\xi\rangle$. ∎
@@
For any $i\in I$ and $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that ${{}_{i}r}(b)=\sum_{b^{\prime}\in\tilde{\mathbf{B}}}\sum_{n\in\mathbb{Z}}d_{b^{\prime},n}^{b}v^{n}b^{\prime}$ for any $b^{\prime}\in\tilde{\mathbf{B}}$ and $n\in\mathbb{Z}$, we have $d_{b^{\prime},n}^{b}\in\mathbb{N}$. Moreover, for any $b^{\prime}\in\tilde{\mathbf{B}}(\xi,\lambda)$ and $n\in\mathbb{Z}$ such that $d_{b^{\prime},n}^{b}\not=0$, we have
@@
Suppose $r(b)=\sum_{b_{1},b_{2}\in\tilde{\mathbf{B}}}c_{b_{1},b_{2}}^{b}b_{1}\otimes b_{2}$. By the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the comultiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, we have $c_{b_{1},b_{2}}^{b}b_{1}\in\mathbb{N}[v,v^{-1}]$. By (13), we have $\sum_{n\in\mathbb{Z}}d_{b^{\prime},n}^{b}v^{n}=c_{\theta_{i},b^{\prime}}^{b}$, and so $d_{b^{\prime},n}^{b}\in\mathbb{N}$. The second statement follows from $\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}(\xi\odot\lambda)$ and [13, Corollary 22.1.5]. ∎
@@
Recall that $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})=\{\pi_{\xi\odot\lambda}(b)\mid b\in\tilde{\mathbf{B}}(\xi,\lambda)\}$, see Definition 4.4. For any $b\in\tilde{\mathbf{B}}(\xi,\lambda)$, by the positivity of the canonical basis $\tilde{\mathbf{B}}$ with respect to the multiplication of $\tilde{\mathbf{f}}$, see Theorem 2.5, we have $\theta_{i}b\in\sum_{b^{\prime}\in\tilde{\mathbf{B}}}\mathbb{N}[v,v^{-1}]b^{\prime}$. By Corollary 4.5, we have
@@
Hence the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ of $\Lambda_{\xi,\lambda}$ has the positivity. Moreover, by Theorem 3.8 and Theorem 4.7, the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ also has the positivity. ∎
@@ 4.4. Generalized case
[…]
@@
[…]
@@
Then we have the corresponding Verma module $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$ and the simple highest weight module $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$ of $\tilde{\mathbf{U}}^{2}$ with the canonical basis $\tilde{\mathbf{B}}^{2}(\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2})$. We restrict them to be $\mathbf{U}$-modules, and can prove that the subspace $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{\mathbf{f}}^{2}$ spanned by $\{x\theta_{\zeta}y\theta_{\lambda}z\mid x,y,z\in\mathbf{f}\}$, where $\theta_{\zeta}=\prod_{i\in I}\theta_{i^{\prime\prime}}^{(\langle i,\zeta\rangle)}\in\tilde{\mathbf{f}}^{2}$, is a $\mathbf{U}$-submodule $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}$. This is analogue to the submodule $\mathbf{f}\theta_{\lambda}\mathbf{f}$ of $\tilde{M}_{\xi\odot\lambda}$. We define the $\mathbf{U}$-module $\Lambda_{\xi,\lambda,\zeta}$ to be the image of $\mathbf{f}\theta_{\zeta}\mathbf{f}\theta_{\lambda}\mathbf{f}$ under the natural projection $\pi_{\xi\odot\lambda\odot\zeta}:\tilde{M}_{\xi\odot\lambda\odot\zeta}^{2}\rightarrow\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2}$. This is analogue to the module $\Lambda_{\xi,\lambda}$. Moreover, we can construct a subset $\tilde{\mathbf{B}}^{2}(\Lambda_{\xi,\lambda,\zeta})\subset\tilde{\mathbf{B}}^{2}(\tilde{\Lambda}_{\xi\odot\lambda\odot\zeta}^{2})$, and prove that it is a basis of $\Lambda_{\xi,\lambda,\zeta}$. This is analogue to the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ of $\Lambda_{\xi,\lambda}$. We can construct a $\mathbf{U}$-module isomorphism $\varphi^{2}:\Lambda_{\xi,\lambda,\zeta}\rightarrow\Lambda_{\xi,\lambda}\otimes\Lambda_{\zeta}$ by using the comultiplication of $\tilde{\mathbf{f}}^{2}$, and prove that the image of $\tilde{\mathbf{B}}^{2}(\Lambda_{\xi,\lambda,\zeta})$ under $\varphi^{2}$ coincides with the canonical basis of the tensor product $\Lambda_{\xi,\lambda}\otimes\Lambda_{\zeta}$. This is analogue to Theorem 3.8 and 4.7.